You have decided to place $175 in equal deposits every month at the beginning of the month into a savings account earning 7.66 percent per year, compounded monthly for the next 14 years. The first deposit is made today. How much money will be in the account at the end of that time period?

Round the answer to two decimal places. Thank you.
Answer #1

r = 7.66%/12 = 0.6383% (monthly), n = 14 * 12 = 168 months

FV = (1 + 0.006383) * 175 * [\frac{(1+0.006383)^{168} - 1}{0.006383}]

FV = 1.006383 * 175 * 299.5992

FV = $52,764.54